OpenStudy (anonymous):

A sample of solid iodine is heated with an electrical coil. If 90.8 Joules of energy are added to a 11.5 gram sample and the final temperature is 39.4oC, what is the initial temperature of the iodine ?

OpenStudy (dtan5457):

Recall that the the formula for heat here is \[q=mc \Delta TC\] q=joules or energy m=mass c=specific heat tc=temperature change So..you plug all that in( the stuff given in question) specific heat of solid iodine=0.427<<according to google 90.8=(11.5)(0.427)(39.4-x)<<as you are heating here, so the initial temperature must be less than 39.4 now you just solve for x.. 90.8=4.9105(39.4-x) distribute like it's algebra 90.8=193.4737-4.9105x -102.6737=-4.9105x x=20.91 (rounded)